OpenStudy (anonymous):

If h(x) = −3x − 10, find h(−3).

OpenStudy (unklerhaukus):

\[h(x) = −3x − 10\] \[h(-3) = −3(-3) − 10\]\[\quad\qquad= ~\dots\]

OpenStudy (anonymous):

im confused

OpenStudy (unklerhaukus):

ok, how are you confused,

OpenStudy (anonymous):

everythings confusing

OpenStudy (unklerhaukus):

all i have done is to substitute x = -3

OpenStudy (unklerhaukus):

to finish we just have to simplify the equation \(h(-3)=-3\times-3-10\)

OpenStudy (unklerhaukus):

does this help \[h(x)=−3(x)−10\]

OpenStudy (anonymous):

otay so i think i got it is he answer h(-3)=-1

OpenStudy (unklerhaukus):

good stuff,

OpenStudy (unklerhaukus):

what is h(0)

OpenStudy (anonymous):

0

OpenStudy (unklerhaukus):

\[h(0)=−3(0) − 10=0-10=-10\]

OpenStudy (unklerhaukus):

when you have bracket in front of a function they indicate the variable

OpenStudy (anonymous):

alright ty i really appericate it